# Data Compression for Fast Online Stochastic Optimization

### v1

@@ 3.4 Performance guarantees
Suppose assumptions 3.3 and 3.4 hold. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (6) at time $t$ satisfies
$-\underline{\psi}_{t}^{K}\leq H_{t}^{K}-H_{t}\leq\bar{\psi}_{t}^{K},$ (12)
@@
$\underline{\psi}_{t}^{K}$ $=\min\left\{\Phi^{K}_{t-1},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\},$
$\bar{\psi}_{t}^{K}$ $=\min\left\{\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\},$
@@
with $\Phi^{K}_{t}$ given in Definition 3.4 and $\Delta_{t}$ given in Lemma 3.2. Furthermore, we obtain the finite-sample probabilistic guarantee
@@
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\underline{\psi}_{t}^{K}\right)\geq 1-\beta_{t-1},$ (13)
where the certificate $H^{K}_{t}+\underline{\psi}_{t}^{K}$ satisfies
@@
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq H_{t}\leq H^{K}_{t}+\underline{\psi}_{t}^{K}\leq H_{t}+\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t-1}.$ (14)
@@
If either the Lipschitz condition or smoothness condition from Assumption 3.4 does not hold, the corresponding term(s) in the bounds are set to $\infty$.
Theorem 3.5 shows that, when the adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$ is constructed with the same radius as nominal DRO, we can relate their optimal values in terms of the distances given in Definition 3.4. We can thus obtain finite-sample guaranteess, with a certificate within a bounded distance from the nominal DRO certificate.
In particular, equation (13) states that with probability $1-\beta_{t-1}$, the out-of-sample performance $\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]$ upper bounds the true optimal value $H_{\star}$, and this value is in turn upper bounded by the in-sample objective $H^{K}_{t}$ and a function $\underline{\psi}_{t}^{K}$ of the clustering discrepancy. This value $H^{K}_{t}+\underline{\psi}_{t}^{K}$ is our certificate.
Using the bounds established in (12), equation (14) then relates this certificate to the one obtained by nominal DRO, providing upper and lower bounds. Specifically, the certificate is at least an upper bound on $H_{t}$, but the increase in suboptimality (as we are solving a minimization problem, a larger certificate is less optimal)is at most $\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}$, which is once again controlled by the clustering discrepancy.
Depending on the smoothness and Lipschitz conditions on the function $f$, we provide multiple bounds for $\underline{\psi}_{t}^{K}$ and $\bar{\psi}_{t}^{K}$. For each application, we choose the best out of the available bounds. Note that in special cases, the terms $\Phi_{t-1}^{K}$ and $\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$ may reduce to 0. Specifically, for DRO problems with support $S={\mbox{\bf R}}^{d}$ (therefore $\Delta_{t}=0$), both values are 0 when the objective function $f$ is affine; we thus recover nominal DRO guarantees regardless of $K$. If the objective function is instead maximum-of-affine, the term involving $\max_{j\leq J}L_{j}$ is reduced to zero. On the other hand, if the objective $f$ is not affine, but we have $J=1$, then $\Phi^{K}_{t-1}=0$, as noted in Definition 3.4.
When the curvature of $f$ is not covered by the special cases above, the clustering discrepancies need to be explicitly computed. When the smoothness condition is satisfied, the term $\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$ can be easily computed using the clustered and non-clustered empirical distributions, and by approximating $\Delta_{t}\simeq 0$. As shown in Lemma 3.2, for ambiguity sets with small radii, this is often the case. When the Lipchitz condition is satisfied, we can compute the bound involving $d^{K}_{t-1,1}$. This is the Wasserstein-1 distance between two empirical distributions, and can be calculated using a linear program. Lastly, the value $\Phi^{K}_{t}$ can be computed regardless of smoothness and Lipschitz conditions, using the dual variables of the solved optimization problem. Therefore, all values in these bounds are readily assessable. In the case of bounded support, we can also use Proposition 3.2 to obtain upper bounds in terms of $\rho$.
Asymptotically, we note that these bounds converge to the following limits.
Theorem 3.6 (Asymptotic guarantee).
@@
Suppose assumptions 3.23.3, and 3.4 hold. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (6) satisfies
@@
where
$\underline{\psi}_{\star}^{K}$ $=\min\left\{\limsup_{t\rightarrow\infty}{\Phi^{K}_{t}},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\},$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\max_{j\leq J}(L_{j}/2)}(D^{K}_{\star,2})^{2},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}.$
@@
This follows from the summability of $\beta_{t}$ and the Borel–Cantelli Lemma [31, Theorem 2.18], as well as Lemma 3.2. The convergence of the bounds follow from the results in Section 3.3. ∎
In order to obtain a closed-form limit for ${\Phi^{K}_{t}}$, we require the convergence of the online algorithm. Below, we prove that the online optimal value converges to a clustering-dependent value, at the same rate of convergence as nominal DRO.
@@
Suppose Assumption 3.2 and the first clause of Assumption 3.4 holds, and let $\varepsilon^{K}_{t-1}\sim O(\varepsilon_{t-1})$, where $\varepsilon_{t-1}$ is computed as in the first clause of Assumption 3.3. With this radius, the optimal value $H^{K}_{t}$ of the compressed DRO problem (6) at time $t\geq\tau$ satisfies
${\mathbf{P}}^{n_{t-1}}\left(H^{K}_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}^{K}_{\star}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t-1},$
and asymptotically, ${\mathbf{P}}^{\infty}$-almost surely we have $H^{K}_{t}\downarrow H^{K}_{\star}$ as $t\to\infty$, where $H^{K}_{\star}$ is the optimal value of the stochastic optimization problem (1) with distribution ${\mathbf{P}}^{K}_{\star}$. In addition, if $f(u,x)$ is lower-semicontinuous in $x$ for all $u\in S$, then any accumulation point $\hat{x}^{K}_{\star}=\lim_{t\to\infty}x_{t}^{K}$ is almost surely, with respect to ${\mathbf{P}}^{\infty}$, an optimal solution to the stochastic optimization problem with distribution ${\mathbf{P}}^{K}_{\star}$.
@@
This follows from Theorem 3.4 and an adaptation of [40, Theorem 3.6], by treating ${\mathbf{P}}^{K}_{\star}$ as the true distribution. ∎
With this result, for special cases, we can obtain a limit for $\Phi^{K}_{t}$.
@@
Suppose the assumptions of Theorem 3.7 hold. When the dual variables $z^{jk}$ of the reformulated problem (8) can be expressed as a continuous function of only the primal variables $x\in\mathcal{X}$, we have $\lim_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star},$ where the latter is defined in Theorem 3.3.
@@
This follows from Theorem 3.7, where we note that the accumulation points $\hat{z}^{jk}_{\star}$ are defined by the accumulation point $\hat{x}^{K}_{\star}$. ∎
For maximum-of-affine $f$ without additional support constraints, the above is always satisfied. Well-known conjugation results show that the variables $z^{jk}=P^{jk}x$, for constant matrices $P^{jk}$ [40, Corollary 5.1].
Under optimal clustering, we obtain the following tighter bounds.
@@
Suppose the assumptions of Theorem 3.6 hold. If the clustering algorithm is optimal with respect to $p=2$ and the $\ell_{2}$-norm, i.e., for all $t$, we achieve the minimum quantization error of $\hat{{\mathbf{P}}}_{t}$ on $K$ clusters, such that $D^{K}_{t,2}=d^{K}_{t,2}=d^{K}_{\star,2}(\hat{{\mathbf{P}}}_{t}),$ then almost surely with respect to ${\mathbf{P}}^{\infty}$,
$\bar{\psi}_{\star}^{K}=\min\left\{{\max_{j\leq J}(L_{j}/2)}(d^{K}_{\star,2}({\mathbf{P}}))^{2},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}.$
On the other hand, suppose the assumptions of Theorem 3.8 hold, and $\|z^{jk}_{t}\|\leq M_{j}$ for $j=1,\dots,J$, $k=1,\dots,K$, and $t\geq 1$. If the clustering-induced coupling is optimal with respect to $p=1$, i.e., $D^{K}_{t,1}=d^{K}_{t,1},$ then almost surely with respect to ${\mathbf{P}}^{\infty}$,
$\underline{\psi}_{\star}^{K}$ $={\Phi^{K}_{\star}}\leq\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}).$
@@
The first result follows from Proposition 3.2 and Theorem 3.2, where the pertinent inequalities are made tight by the assumption of optimal clustering. For the second result, note that $\Phi^{K}_{t}$ is computed with respect to the clustering-induced coupling, i.e., each $\hat{u}$ is associated with the centroid $\bar{u}^{k}_{t}$ of the cluster it belongs to. If this is equivalent to the optimal coupling, then by the Cauchy-Schwarz inequality,
@@
which holds for all $t$, $j$, and $\hat{u}$. This implies
$\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|=\max_{j\leq J}M_{j}d^{K}_{t,1}.$
The result follows by taking the limit $t\to\infty$. ∎
Note that, the requirement on $z^{jk}_{t}$ is once again satisfied by maximum-of-affine $f$ without extra support constraints. Furthermore, while we only prove that $\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}d^{K}_{t,1}$ under the conditions above, the large gain from the Cauchy-Schwarz inequality generally implies that this relationship, even without optimal coupling.
In all cases, our bounds scale with the clustering quality; when the clustering algorithm is optimized, the bound is tighter. Asymptotically, the performance-gap between the online solution and the nominal DRO solution is controlled by different distance metrics between the true distribution and its discretization to $K$ atoms, and, in the case of Lipschitz $f$, is upper bounded by the Wasserstein-1 distance. Theorem 3.7 shows a direct correlation between the optimality of the clustering procedure and the optimality of the online solution; the algorithm will converge to some value dependent on the supports of the final clustering, so it follows that a well-chosen clustering procedure will lead to a better performing solution.
For all these bounds, we can verify using either Proposition 3.2 or Theorem 3.2 that as the number of clusters $K\to\infty$, we recover the DRO asymptotic performance guarantee: as ${t\to\infty}$, we have ${\mathbf{P}}^{\infty}$-almost surely $H^{K}_{t}\rightarrow H_{\star}$. This illustrates the importance of finding a $K<\infty$ which balances the tradeoff between performance and computational effort. In fact, this leads to an interesting result: if we do not have a strict budget $K$, but would like to decrease computational effort, we can obtain the following more controllable bounds.
@@
Suppose assumptions 3.3 and 3.4 hold. If we fix a radius $\eta$ such that all data-points are either clustered into an existing ball with at most radius $\eta$, or initialized as a new cluster, then ${\mathbf{P}}^{\infty}$-almost surely,

### v2

@@ 3.3.1 Concave case
Suppose Assumptions 2.12.2, and 3.2 hold, $f$ is concave in $u$, and $S={\bf R}^{d}$. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t$ satisfies
$H_{t}\leq H_{t}^{K}\leq H_{t}+\bar{\psi}_{t}^{K},\quad\bar{\psi}_{t}^{K}=\min\left\{{\frac{L}{2}}(D^{K}_{t,2})^{2},M(2\varepsilon_{t}+W^{K}_{t,1})\right\}$ (12)
where $H_{t}$ is the nominal DRO value. Furthermore, we obtain the finite-sample probabilistic guarantee
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{t}}\right)\geq 1-\beta_{t},$ (13)
where the certificate $H^{K}_{t}+\rho_{t}$ satisfies
${\mathbf{P}}^{{t}}\left(H_{\star}\leq H_{t}+\rho_{t}\leq H^{K}_{t}+\rho_{t}\leq H_{t}+\rho_{t}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t}.$ (14)
If any among the Lipschitz and smoothness conditions from Assumption 2.1 does not hold, the corresponding term in $\bar{\psi}_{t}^{K}$ is set to $\infty$.
Theorem 3.4 shows that, when the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is constructed with the same radius as nominal DRO, we can relate their optimal values in terms of this radius and the distances given in Definition 3.4. This yields a finite-sample guarantee, with a certificate within a bounded distance from the nominal DRO certificate.
In particular, equation (13) states that with probability $1-\beta_{t}$, the out-of-sample performance $\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]$ upper bounds the true optimal value $H_{\star}$, and this value is in turn upper bounded by the in-sample objective $H^{K}_{t}$ and the nominal DRO residual $\rho_{t}$, which depends on the radius $\varepsilon_{t}$. Using the bounds established in (12), equation (14) relates this certificate to the one obtained by nominal DRO, providing upper and lower bounds. Specifically, the certificate is an upper bound on the nominal certificate $H_{t}+\rho_{t}$, but the increase in suboptimality (as we are solving a minimization problem, a larger certificate is less optimal) is at most $\bar{\psi}_{t}^{K}$. Similar to the nominal residual, which is the price paid in exchange for smaller radii, $\bar{\psi}_{t}^{K}$ is the price we pay for increased computational efficiency. As the radius $\varepsilon_{t}$ decreases over time at the chosen rate, which could be as fast as $\tilde{O}(n_{t}^{-1/2})$, this price converges to, or already is, a function of only the clustering distance. Note that the distances for the clustering discrepancy are defined with orders $p=1$ or $p=2$, unrelated to the order $r$ of the ambiguity set. The order $r$ instead affects the bounds through the radius $\varepsilon_{t}$ and the residual $\rho_{t}$.
Depending on the smoothness and Lipschitz conditions on the function $f$, we can choose the best out of the two clauses given for $\bar{\psi}_{t}^{K}$. Note that when $f$ is affine, $\bar{\psi}_{t}^{K}$ reduces to 0, and we recover nominal DRO guarantees regardless of $K$. Otherwise, when the Lipchitz condition is satisfied, we can compute the bound involving $W^{K}_{t,1}$. This is the Wasserstein-1 distance between two empirical distributions, and can be calculated using a linear program. When the smoothness condition is satisfied, $D_{t,2}^{K}$ can be easily computed using the empirical distributions.
Asymptotically, if $\varepsilon_{t}$ is chosen through Theorem 2.1, the bound converges as in the following theorem. Note that while $\varepsilon_{t}$ chosen through Theorem 2.2 allows smaller finite-sample differences between the certificates, i.e., $\tilde{O}(n_{t}^{-1/2})$ scales better than $O(n_{t}^{-p/d})$, in the asymptotic case we can assume $O(n_{t}^{-p/d})$ to converge to 0 for any $d$. In fact, our method is more compatible with asymptotic analysis than nominal DRO, as our computational complexity is fixed regardless of the sample size.
Theorem 3.5 (Concave asymptotic guarantee).
Suppose Assumptions 2.12.23.1 and 3.2 hold, $f$ is concave, $S={\bf R}^{d}$, and $\varepsilon_{t},\beta_{t}$ are constructed according to Theorem 2.1 for all $t$. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\ \leq\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+\bar{\psi}_{\star}^{K}\right\}=1,$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\frac{L}{2}}(D^{K}_{\star,2})^{2},M{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\}.$
This shows that in the asymptotic case, the performance gap between the online solution and the nominal DRO solution is controlled by our defined distances between the true distribution and a chosen discretization to $K$ atoms. It follows that a well-chosen clustering algorithm and a balanced $K$ are crucial for good performance.
3.3.2 General case
We now extend these results to the maximum-of-concave case with arbitrary support $S$, which requires the following lemma. The proofs of this lemma and the upcoming results are in Appendix A.1.
Lemma 3.2 (Effect of the support).
Let $\Delta_{t}=\tilde{H}_{t}-H_{t}$, where $H_{t}$ is the value of the DRO problem with arbitrary support $S$, while $\tilde{H}_{t}$ fixes $S={\bf R}^{d}$, i.e., $\tilde{H}_{t}=\max_{\mathbf{Q}\in\hat{\mathcal{P}}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$, $\hat{\mathcal{P}}_{t}=\{\mathbf{Q}\in\Xi_{r}({{\bf R}^{d}})\mid W_{r}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\}.$ If the sequence of $\varepsilon_{t}$ is constructed according to Theorem 2.1, we have ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Delta_{t}=0\right\}=1.$ For any $\varepsilon_{t}\leq 1$, if $f$ satisfies clause 1 or 3 of Assumption 2.1, with $r\geq 2$ in the latter case, then $\Delta_{t}=O(\varepsilon_{t})$.
This lemma formalizes the discrepancy between nominal DRO objectives with and without support constraints. This discrepancy appears in the following bounds.
Theorem 3.6 (General finite sample guarantee).
Suppose Assumptions 2.12.2, and 3.2 hold. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t$ satisfies
$\underline{\psi}_{t}^{K}\leq H_{t}^{K}-H_{t}\leq\bar{\psi}_{t}^{K},$ (15)
@@ 3.3.2 General case
$\underline{\psi}_{t}^{K}$ $=\min\left\{\Phi^{K}_{t},\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1}),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f_{x^{K}_{t}}\|_{L_{2}(\hat{{\mathbf{P}}}^{K}_{t})}(2\varepsilon_{t}+W^{K}_{t,2})+\frac{L}{2}(2\varepsilon_{t}+W^{K}_{t,2})^{2}}\right\},$
$\bar{\psi}_{t}^{K}$ $=\min\left\{\Delta_{t}+{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{t,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1})\right\},$
@@
with $\|\nabla f_{x}\|_{L_{2}(\mathbf{Q})}=(\int\|\nabla f(u,x)\|^{2}d\mathbf{Q})^{1/2},$ $\Phi^{K}_{t}$ given in Definition 3.4, and $\Delta_{t}$ given in Lemma 3.2. Furthermore, we obtain the finite-sample probabilistic guarantee
@@
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{t}}+\underline{\psi}_{t}^{K}\right)\geq 1-\beta_{t},$ (16)
where the certificate $H^{K}_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}$ satisfies
@@
${\mathbf{P}}^{{t}}\left(H_{\star}\leq H_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}\leq H^{K}_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}\leq H_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t}.$ (17)
@@
If any among the Lipschitz and smoothness conditions from Assumption 2.1 does not hold, the corresponding term(s) in the bounds are set to $\infty$.
Theorem 3.6 shows that, for maximum-of-concave $f$ with arbitrary support, the online objective $H^{K}_{t}$ is no longer guaranteed to be an upper bound on $H_{t}$. This leads to performance guarantees and performance gaps with additional terms compared to that of Theorem 3.4. However, the additional terms are also functions of the ambiguity set radius $\varepsilon_{t}$ and the clustering distances, and can be readily approximated or computed. For instance, $\Delta_{t}\approx O(\varepsilon_{t})$, and the value $\Phi^{K}_{t}$ can be computed using the dual variables of the solved optimization problem. Asymptotically, we can also obtain limits for the generalized bounds.
Theorem 3.7 (General asymptotic guarantee).
@@
Suppose Assumptions 2.12.23.1 and 3.2 hold, and $\varepsilon_{t},\beta_{t}$ are constructed according to Theorem 2.1 for all $t$. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
@@
$\underline{\psi}_{\star}^{K}$ $=\min\left\{\limsup_{t\rightarrow\infty}{\Phi^{K}_{t}},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f_{\hat{x}^{K}_{\star}}\|_{L_{2}({\mathbf{P}}^{K}_{\star})}W^{K}_{\star,2}+\frac{L}{2}(W^{K}_{\star,2})^{2}}\right\},$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{\star,2})^{2},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\},$
where $\hat{x}^{K}_{\star}$ is the accumulation point of $x^{K}_{t}$ that yields the largest value of the bound.
@@
This follows from the summability of $\beta_{t}$ and the Borel–Cantelli Lemma [28, Theorem 2.18], as well as Lemma 3.2. The convergence of the bounds follow from the results in Section 3.2. The existence of $\hat{x}^{K}_{\star}$ follows from the compactness of $\mathcal{X}$.
In order to obtain a closed-form limit superior for $\Phi^{K}_{t}$}, we require the convergence of the online algorithm. Below, we prove that the online optimal value converges to a clustering-dependent value, at the same rate of convergence as nominal DRO.
@@
Suppose Assumption 3.1 and the Lipschitz clause of Assumption 2.1 holds, and let $\varepsilon^{K}_{t}\sim O(\varepsilon_{t})$, where $\varepsilon_{t}$ is computed according to Theorem 2.1. With this radius, the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t\geq\tau$ satisfies
${\mathbf{P}}^{{t}}\left(H^{K}_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}^{K}_{\star}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t},$
[…]
@@
This follows from Theorem 3.3 and an adaptation of [38, Theorem 3.6], by treating ${\mathbf{P}}^{K}_{\star}$ as the true distribution. ∎
With this result, for special cases, we can obtain a limit superior for $\Phi^{K}_{t}$.
@@
Suppose the assumptions of Theorem 3.8 hold. When the function $f(u,x)$ is maximum-of-affine in both $u$ and $x$, and the support $S={\bf R}^{d}$, we have $\limsup_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star},$ where the latter is defined in Theorem 3.2.
@@
When $f$ is maximum-of-affine and $S={\bf R}^{d}$, well known conjugation results show that the variables $z^{jk}=P^{jk}x$, for constant matrices $P^{jk}$ [38, Corollary 5.1]. It follows that the accumulation points $\hat{z}^{jk}_{\star}$ exist as continuous functions of an accumulation point $\hat{x}^{K}_{\star}$, which exist by Theorem 3.8.
The above results hold in general, for any clustering algorithm satisfying Assumption 3.1. For completeness, below we instead assume optimal clustering, and state the best bounds we can obtain.
@@
Suppose Assumptions 2.12.23.1 and 3.2 hold. If the clustering algorithm is optimal with respect to $p=2$ and the $\ell_{2}$-norm, i.e., for all $t$, we achieve the minimum quantization error of $\hat{{\mathbf{P}}}_{t}$ on $K$ clusters, such that $D^{K}_{t,2}=W^{K}_{t,2}=W^{K}_{\star,2}(\hat{{\mathbf{P}}}_{t}),$ then almost surely with respect to $\mathbf{P}$$^{\infty}$,
$\bar{\psi}_{\star}^{K}=\min\left\{{\max_{j\leq J}\frac{L_{j}}{2}}(W^{K}_{\star,2}({\mathbf{P}}))^{2},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\}.$
When $f(u,x)$ is maximum-of-affine in both $u$ and $x$, and the support $S={\bf R}^{d}$, if the clustering-induced coupling is optimal with respect to $p=1$, i.e., $D^{K}_{t,1}=W^{K}_{t,1},$ then almost surely with respect to $\mathbf{P}$$^{\infty}$,
${\underline{\psi}_{\star}^{K}\leq\Phi^{K}_{\star}}\leq\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}.$
@@
The first result follows from Proposition 3.1 and Theorem 3.1, where the pertinent inequalities are made tight by the assumption of optimal clustering. For the second result, note that $\Phi^{K}_{t}$ is computed with respect to the clustering-induced coupling, i.e., each $\hat{u}$ is associated with the mean $\bar{u}^{k}_{t}$ of the cluster it belongs to. If this is equivalent to the optimal coupling, then by the Cauchy-Schwarz inequality,
@@
where $\|z^{jk}_{t}\|\leq M_{j}$ $\forall j\leq J$, $k\leq K$, and $t\geq 1$ by the conjugates of maximum-of-affine functions and the compactness of $\mathcal{X}$. This holds for all $t$, $j$, and $\hat{u}$, which implies
$\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|=\max_{j\leq J}M_{j}W^{K}_{t,1}.$
The result follows by taking the limit superior as $t\to\infty$.
In all cases, our bounds scale with the clustering quality; when the clustering algorithm is optimized, the bound is tighter. Theorem 3.8 shows a direct correlation between the optimality of the clustering procedure and the optimality of the online solution; the algorithm will converge to a value dependent on the supports of the final clustering, so it follows that a well-chosen clustering procedure will lead to a better performing solution. For all these bounds, we can verify using either Proposition 3.1 or Theorem 3.1 that as the number of clusters $K\to\infty$, we recover the DRO performance guarantees. This illustrates the importance of finding a $K<\infty$ which balances the tradeoff between performance and computational effort.
In fact, this leads to an interesting result: if we do not have a strict budget $K$, but would like to decrease computational effort, we can obtain the following more controllable bounds.
@@
[…]